# VersaTune: An Efficient Data Composition Framework for Training Multi-Capability LLMs

### v4

@@ 5.3. Results
VersaTune is efficient across different models in both scenarios. Table 4 and Figure 5 show that VersaTune consistently outperforms other baseline methods across different models in terms of domain-specific capabilities. Compared to the uniform distribution of data across domains, VersaTune enhances downstream task performances by , which further underscores the effectiveness of our data composition strategy for enhancing the model’s overall multi-domain capabilities during the supervised fine-tuning phase. Since we have not conducted domain-specific refinement for domains outside the current five specific domains, the models’ performance gains on general benchmarks are not as noticeable. For domain expansion scenarios, VersaTune has nearly maintained training efficiency while reducing the model’s loss of competencies in other domains by comparing to 100% specific domain fine-tuning, as depicted in Table 5 and Figure 6, where we averaged the experimental results from Baichuan-2-7B and Baichuan-2-13B. Three interesting phenomena are observed from the outcomes:
@@
Absolute Count vs. Proportion. Notably, by the second epoch of training, the degradation in performance across non-target domains tends to be mitigated to some extent, and in some cases, there is even a positive trend in capability enhancement. We attribute this phenomenon to the fact that the absolute number of samples for each domain, relative to the distribution of domain samples, has a predominant influence on model performance at this stage.
@@
Domain Interactions. Domains are not entirely orthogonal to each other, and there is a degree of mutual reinforcement among them. For instance, enhancing capabilities in the medicine domain can, to a certain degree, boost performance in the science domain, as illustrated in Figure 6.
@@
Domain Mastery Efficiency. From the slope of the target domain performance increase in Figure 6 (b), it is evident that the model’s efficiency in mastering knowledge of a specific domain diminishes over training. In other words, as training progresses, the model’s grasp of the current target domain approaches saturation, while its performance on non-target domains declines sharply. Consequently, greater emphasis should be placed on mitigating losses in non-target domains during this phase, aiming to strike a balance between domain expansion and the salvage of capabilities in non-target domains, which is also demonstrated in Figure 7.
Knowledge consistency training boosts performance. In Table 4 and Figure 5, we also present the experimental results of data composition strategies that allocate domain data in a manner inversely proportional to the pre-existing knowledge distribution detected within each domain. As expected, the inverse distribution strategy yielded even lower performance compared to the simplest approach of uniform distribution, which evenly distributes data across all domains. This finding underscores the importance of aligning domain data ratios with the inherent knowledge distribution of the model during training, which proves the efficacy of knowledge consistency training stated in Section 3.1.1.
Dynamic adjustment enhances the robustness. During the process of cultivating multiple capabilities, we compared VersaTune with fixed domain weights referring to the knowledge distribution obtained from probing the target model prior to supervised fine-tuning, namely VersaTune Constant, to ablate the component of dynamic adaptation in Algorithm 2. Table 4 and Figure 5 demonstrated the high robustness of VersaTune, which dynamically adjusts domain weights throughout the training process by continuously monitoring the learnable potential within each domain. In contrast, training with fixed domain weights exhibits certain fluctuations. A key reason for this phenomenon is the distribution of domain knowledge mastered by the model changes during training, and the learning efficiency varies among domains. Therefore, dynamically adjusting domain data weights based on the model’s feedback at different stages of training is crucial.
@@
Establishing proportion thresholds for specific domains counts during domain expansion. We consider conducting a comparative analysis between the outcomes of VersaTune and those implementing an unconditional dynamic increase of the specific domain, where we remove the implementation of Line 8 in Algorithm 3, to ablate the component of criteria for determining the upper limit of domain expansion in Algorithm 3. In Figure 7, we present the trends in the overall multi-domain performance of the models under specific domain expansion for each domain. It can be observed that, for the majority of domains, the gap in average multi-task performance between models trained with VersaTune and those without an upper limit on domain proportion becomes increasingly pronounced after the second or third epoch. We speculate that it is because by the later stages of supervised fine-tuning, models’ proficiency in the target domain approaches saturation. Further increasing the proportion of the current domain provides diminishing returns and can lead to a significant loss of performance in other domains. Notably, between the second and third epochs of fine-tuning, the model reaches a balance where the efficiency of improvement in the target domain is matched by the rate of performance degradation in non-target domains. The finding shows that the criteria for determining the upper limit on the proportion of a specific domain during domain expansion, has mitigated the loss of capabilities in other domains experienced by the target model during the fine-tuning process. Concurrently, it ensures gains in the capacity for the current domain of interest.
6. Related Work
Data Reweighting for LLM Training. Data reweighting maintains full access to the entire dataset while adjusts the relative importance of each instance for various target tasks, which is essential for both pretraining and fine-tuning stages of LLMs (Wang et al., 2023). During the pretraining stage, DoReMi (Xie et al., 2024b) and DoGE (Fan et al., 2023) employ lightweight proxy models to estimate weights for different data sources, which are subsequently applied to the formal training of LLMs. Furthermore, Sheared LLaMA (Xia et al., 2023) implements an online variant of DoReMi, where the loss reference provided by the proxy model is replaced with the state of the pre-trained model from the previous training step. Additionally, ODM (Albalak et al., 2023) leverages the differences in loss or perplexity across domains within a multi-armed bandit framework to dynamically adjust the data distribution along training steps. As for the SFT phase, Dong et al. (Dong et al., 2023) focus on enhancing the model’s math reasoning, coding, and general human-aligning abilities through a dual-stage mixed fine-tuning strategy. However, the mixing ratios for different domains rely heavily on enumeration and empirical methods, and the covered domains are not holistic. This paper provides a comprehensive overview of the model’s capabilities across various domains during the SFT stage and proposes appropriate multi-ability fine-tuning methods.
Knowledge Detection in LLMs. Investigating the knowledge contained in current LLMs is essential for guiding their subsequent training (Chang et al., 2024). The knowledge encompasses multiple dimensions, such as different domain sources and task attributes. Existing work in LLM knowledge detection primarily focuses on prompting and calibration. Directly prompting the model to generate sequences and extracting confidence scores from the model (Gekhman et al., 2024; Kadavath et al., 2022; Kuhn et al., 2023; Manakul et al., 2023) is a common strategy. However, such approaches are highly dependent on prompt design and task selection, which can introduce bias into the assessment. Other studies have attempted to infer the training data mixtures used in previous training stages of LLMs (Antoniades et al., 2024; Hayase et al., 2024; Hu et al., 2022; Ye et al., 2022). The essence of these studies is to evaluate the current knowledge state of the models and provide targeted strategies for data organization and management in subsequent training phases.
7. Conclusion
Large-scale pretrained models, such as Large Language Models (LLMs), have demonstrated exceptional performance across multiple domains due to their emergent properties. Versatile capability across multiple domains is of paramount importance, which is typically cultivated during the Supervised Fine-Tuning (SFT) phase of LLMs. However, existing data composition methods for the SFT stage of LLMs often focus on single domain tasks, leading to significant degradation in performance across other domains. Therefore, there is an urgent need for an effective data composition strategy during the SFT phase. In this paper, we introduce VersaTune, a data composition framework designed to enhance the multi-domain capabilities of models during the supervised fine-tuning phase, which is based on the domain knowledge distribution of the target model. Experimental results from different models and downstream tasks across various domains demonstrate that VersaTune achieves excellent training outcomes in both scenarios of overall multi-domain enhancement and flexible expansion of specific domains.

### v5

@@ D.2 Flexible Domain Expansion
D.2.1 Performance Variations in Target and Non-Target Domains
Here we exhibit the performance of the target domain and other non-target domains under the domain expansion scenario, as realized by Algorithm 3. Figure 3, Figure 10-13 illustrate the changes in target domain capabilities and non-target domain capabilities during the fine-tuning process when focused on a specific domain, providing experimental results for flexible domain expansion. In each figure, the stacked group bar chart (left) depicts the percentage change in performance for non-target domains relative to their pre-fine-tuning states, with the positive direction on the y-axis indicating performance improvement and the negative direction signifying a decline. The line chart (right) represents the overall change across all non-target domains for three distinct strategies, with color legends corresponding to those of the line chart on the right. The right-side chart depicts the percentage increase in performance for the current target domain. We employed the Qwen-2.5-7B and Qwen-2.5-14B models, and the mean percentage change in model performance when focusing on domain enhancement is presented in both the stacked group bar chart and the line chart. Three interesting phenomena are observed from the outcomes:
@@ D.2.1 Performance Variations in Target and Non-Target Domains
Absolute Count vs. Proportion. Notably, by the second epoch of training, the performance degradation across non-target domains tends to be mitigated to some extent, and there is even a positive trend in capability enhancement in some cases. We attribute this phenomenon to the fact that the absolute quantity of instances for each domain, relative to the domain distribution, has a predominant influence on model performance at this stage.
@@
Domain Interactions. Domains are not entirely orthogonal to each other, and there is a degree of mutual reinforcement among them: (1) Enhancing capabilities in the medicine domain can boost performance in the science domain to a certain degree (Figure 3). (2) Models’ capabilities in law and finance are mutually reinforcing, promoting each other’s development (Figure 10 and Figure 11). (3) Augmenting the model’s code-related capabilities can also, to some extent, improve its ability to solve scientific problems, which is likely due to the shared reasoning and logical structuring required across these domains (Figure 13).
@@
Domain Mastery Efficiency. From the slope of the target domain performance increase in Figure 3, Figure 10-13 (b), it is evident that the model’s efficiency in mastering knowledge of a specific domain diminishes over training. In other words, as training progresses, the model’s grasp of the target domain approaches saturation, while its performance on non-target domains declines sharply. Consequently, greater emphasis should be placed on mitigating losses in non-target domains during this phase, aiming to strike a balance between domain expansion and the salvage of capabilities in non-target domains, which is also shown in Figure 14.
D.2.2 Importance of Proportion Thresholds
@@
Here we describe the significance of establishing proportion thresholds for specific domains during domain expansion in detail. We compare VersaTune with those implementing an unconditional dynamic increase of the specific domain, where we remove the implementation of Line 8 in Algorithm 3, to ablate the component of criteria for determining the upper limit of domain expansion. In Figure 14, we present the trends in the overall multi-domain performance of the models under specific domain expansion for each domain. It can be observed that, for the majority of domains, the gap in average multi-task performance between models trained with VersaTune and those without an upper limit on domain proportion becomes increasingly pronounced after the second or third epoch. We deduce that this occurs due to the fact that as training progresses, the models’ ability to learn within the target domain becomes nearly maximized. Enhancing the emphasis on the current domain beyond this point yields marginal benefits and may even result in a substantial degradation of performance in other domains. Notably, between the second and third epochs of supervised fine-tuning, the model reaches a balance where the efficiency of improvement in the target domain is matched by the rate of performance degradation in non-target domains. The finding shows that the criteria for determining the upper limit on the proportion of a specific domain during domain expansion, has mitigated the loss of capabilities in other domains experienced by the target model during the fine-tuning process. Moreover, it ensures gains in the capacity for the current domain of interest.
In summary, VersaTune exhibits the following properties:
Efficient. VersaTune employs distribution consistency training of the domain knowledge proportion during models’ SFT stage, providing an efficient data composition strategy for enhancing versatile capabilities (for C2).
Appendix E Prompts
We present the prompts that are employed throughout our pipeline in VersaTune . Only the English version is presented due to LaTeX compilation issues with non-English languages.